OpenStudy (anonymous):

Find the number which completes the square. x^2+24x=c

OpenStudy (campbell_st):

have the b value then square it (24/2)^2 = 144 since the quadratic is x^2 + 24x - c= 0 to make it a perfect square c = -144 - x - gives a positive

OpenStudy (apoorvk):

you know, \[(a+b)^2 = a^2 + 2ab + b^2\] now comparing with above equation, a=1x 2ab=24x so, b=12 and ofcourse, b^2 is the"-c" here (minus because you bring it to the LHS later) so c= -144